### v4

# GOAT-Bench: Safety Insights to Large Multimodal Models through Meme-Based Social Abuse

Abstract
The exponential growth of social media has profoundly transformed how information is created, disseminated, and absorbed, exceeding any precedent in the digital age. Regrettably, this explosion has also spawned a significant increase in the online abuse of memes. Evaluating the negative impact of memes is notably challenging, owing to their often subtle and implicit meanings, which are not directly conveyed through the overt text and image. In light of this, large multimodal models (LMMs) have emerged as a focal point of interest due to their remarkable capabilities in handling diverse multimodal tasks. In response to this development, our paper aims to thoroughly examine the capacity of various LMMs (e.g., GPT-4o) to discern and respond to the nuanced aspects of social abuse manifested in memes. We introduce the comprehensive meme benchmark, GOAT-Bench, comprising over 6K varied memes encapsulating themes such as implicit hate speech, sexism, and cyberbullying, etc. Utilizing GOAT-Bench, we delve into the ability of LMMs to accurately assess hatefulness, misogyny, offensiveness, sarcasm, and harmful content. Our extensive experiments across a range of LMMs reveal that current models still exhibit a deficiency in safety awareness, showing insensitivity to various forms of implicit abuse. We posit that this shortfall represents a critical impediment to the realization of safe artificial intelligence. The GOAT-Bench and accompanying resources are publicly accessible at https://goatlmm.github.io/, contributing to ongoing research in this vital field.
Disclaimer: This paper contains content that may be disturbing to some readers.
Figure 1: Performance on our GOAT-Bench of a broad range of representative LMMs, like CogVLM (Wang et al., 2023a), InstructBLIP (Dai et al., 2023), LLaVA-1.5 (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Qwen-VL (Bai et al., 2023), and GPT-4V(ision) (OpenAI, 2023). GPT-4V achieves the best overall performance from five different perspectives.
1 Introduction
Remarkable progress has been made in the field of natural language processing (NLP), marked by the emergence of large language models (LLMs) equipped with billions of parameters and pre-trained on extensive data (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, b). Through instruction fine-tuning and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), cutting-edge LLMs, like ChatGPT and GPT-4 (OpenAI, 2023), have demonstrated exceptional proficiency in comprehending and executing human instructions, even without tuning the parameters to offer powerful zero-shot or few-shot performance (Wei et al., 2021; Kojima et al., 2022; Lin et al., 2022; Wei et al., 2023). In a related vein, efforts have been directed towards equipping large multimodal models (LMMs) (Team et al., 2023; OpenAI, 2023; Zhu et al., 2023; Liu et al., 2023a; Gong et al., 2023; Dai et al., 2023) with similar capabilities, extending their proficiency to a range of multimodal tasks (Hudson and Manning, 2019; Lu et al., 2022; Fu et al., 2023). This dual advancement in both unimodal and multimodal models marks a significant stride in the field of artificial intelligence, paving the way for more integrated and versatile applications.
Recent literature (Huang et al., 2023) has highlighted that LLMs can achieve remarkable performance in addressing the social abuse inherent in natural languages (An et al., 2021). This area of application is notably more ‘subjective’ compared to other NLP tasks on social media (Lin et al., 2021) and presents significant challenges, even for human evaluators, who often demonstrate lower agreement rates. In addition to language expression, the proliferation of social media has led to the rise of a distinctive multimodal phenomenon: memes (Zenner and Geeraerts, 2018). A typical meme consists of an image coupled with a brief textual component (Pramanick et al., 2021a; Lin et al., 2023). Their ease of sharing enables memes to rapidly disseminate across various online platforms. While commonly perceived with humor, memes are increasingly being co-opted for online abuse, utilizing a strategic interplay of text and image that often relates to current political and socio-cultural events (Sharma et al., 2022; Lin et al., 2024). This evolving use of memes highlights their potential impact on the digital landscape, especially in the context of contemporary discourse. The challenge in detecting social abuse in memes lies in the subtlety of their content, where the underlying meanings are not immediately apparent in the text and images of the memes (Pramanick et al., 2021b). This task demands extensive commonsense knowledge and reasoning skills, built upon a precise interpretation of both the textual and visual elements of memes.
However, in contrast to LLMs, the abilities and constraints of LMMs in handling relatively subjective and nuanced tasks, particularly in identifying online abusive memes within multimodal inputs, are yet to be fully explored (Awadalla et al., 2023; Chen et al., 2023). These tasks are integral to understanding social phenomena, demanding nuanced social judgment and decision-making skills. Given that LMMs are trained on extensive and diverse image-text corpora, showcasing impressive generalization skills (Liu et al., 2023a; Chen et al., 2024), it becomes imperative to assess both the potential strengths and challenges they may face in addressing meme-based social abuse. This line of inquiry is especially pertinent to safety insights, focusing on how LMMs discern and process the intricate combination of visual and textual elements in a manner that is both accurate and socially conscientious so that its output does not cause harm or discomfort.
To this end, we introduce the GOAT benchmark, i.e., GOAT-Bench111 Goat is often depicted as evil in memes due to its historical and mythological associations with devilish figures, reflected in its unique features like horns and eyes., a comprehensive and specialized dataset designed to evaluate large multimodal models through meme-based multimodal social abuse. GOAT-Bench comprises over 6K diverse memes, encompassing a range of themes including hate speech and offensive content. Our focus is to assess the ability of LMMs to accurately identify online abuse, specifically in terms of hatefulness, misogyny, offensiveness, sarcasm, and harmfulness. We meticulously control for the granularity of each specific meme task to facilitate a detailed analysis. Furthermore, we extend our evaluation to assess the effectiveness of thought chains, few-shot demonstrations, and supervised fine-tuning in discerning the underlying implications of memes for deducing their potential threat to online safety.
Leveraging the GOAT-Bench, we assess the capability of various LMMs as exemplified in Figure 1, to detect meme-based social abuse through complex reasoning. We conceive of GOAT-Bench as a comprehensive suite of benchmarks, with subsets tailored to evaluate specific model abilities and analyze the key strengths and weaknesses of LMMs. For instance, in tasks focused on sexism, such as misogyny, we observe significant disparities among LMMs; in broader tasks like harmfulness, which demand advanced background knowledge and reasoning, the performance of LMMs tends to be more concentrated and generally modest. Our aim with GOAT-Bench is to enable researchers to gain a multi-dimensional understanding of their models’ abilities in addressing meme-based social abuse and to contribute to the advancement of safety insights in LMMs, as a critical step in preventing the escalation of online social abuse and ensuring the stability and cohesion of diverse communities.
Figure 2: GOAT-Bench is a comprehensive dataset that tackles the five interwoven meme tasks.
Our contributions are summarized as follows in three folds:
  • We introduce GOAT-Bench, the first advanced comprehensive testbed consisting of 6,626 memes, spanning five tasks of varied complexity related to social abuse, tailored to systematically dissect and analyze the safety insights of LMMs.
  • We extensively evaluate the performance of the 11 most cutting-edge large multimodal models on the GOAT-Bench benchmark. Our findings indicate that GOAT-Bench poses significant challenges to the current state-of-the-art models. Remarkably, the top-performing GPT-4V only attains an overall macro-averaged F1 score of 70.29%, with all the other open-source LMMs registering scores below 62%.
  • Our study conducts an empirical and exhaustive analysis of the performance disparities among various closed-source and open-source LMMs on GOAT-Bench, employing diverse deployment strategies such as chain-of-thought, in-context learning prompts, and the instruction-tuning paradigm. It is our aspiration that our research, along with the GOAT-Bench benchmark, lays the groundwork for future advancements in this realm, catalyzing the evolution of increasingly sophisticated and proficient large multimodal models.
2 The GOAT Benchmark
2.1 Tasks
With the objective of methodically assessing the multimodal safety and associated reasoning capabilities of LMMs in relation to meme-based social abuse, we formulate five research questions, as shown in Figure 2. Subsequently, we curate a collection of abusive memes from diverse sources to address these questions, as summarized in Table 1.
Task I: Hatefulness. Hateful memes pose a direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease (Kiela et al., 2020), where the attack is defined as violent or dehumanizing (comparing people to non-human things, e.g. animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hate speech in memes. This raises the question of whether LMMs can reason over the implicit meaning behind the memes to distinguish hatefulness. To investigate this issue, we have selected meme samples from the FHM dataset (Kiela et al., 2020), which was released by Facebook as part of a challenge to crowd-source multimodal harmful meme detection in hate speech solutions. To assess the performance of multimodal models in handling synthetic memes with added confounders, we have organized the unseen test samples where unimodal information is insufficient for detection and deep multimodal reasoning is required. The data contains hateful memes targeting various vulnerable groups in categories including Religion, Race, Gender, Nationality, and Disability.
Task II: Misogyny. A meme is misogynous if it conceptually describes a sexist or hateful scene (weak or strong, implicitly or explicitly) having as target a woman or a group of women, in the form of shaming, stereotype, objectification and/or violence (Fersini et al., 2022). This prompts the inquiry into whether LMMs possess the capability to effectively snippy niche abusive content that specifically targets certain groups, such as women. To dive into this issue, we resorted to meme samples from the MAMI dataset (Fersini et al., 2022) to reflect the detection capability of LMMs towards social abuse for female victims.
Task III: Offensiveness. Offensiveness spreads an idea or emotion that intends to damage the social identity of the target person, community, or lower their prestige (Suryawanshi et al., 2020a). In contrast to hateful memes, which target specific groups with expressions of animosity or discrimination, offensive memes encompass a wider spectrum of language or image that has the potential to offend, hurt, or provoke anger among individuals. They may not be directed towards a particular group and can affect individuals in a general sense (Drakett et al., 2018). In light of this understanding, it becomes imperative to investigate the efficacy of LMMs in scrutinizing online multimodal offensive meme content, known for its intent to provoke distress or embarrassment through its use of rudeness. To this end, we organized the offensive memes by utilizing the MultiOFF dataset (Suryawanshi et al., 2020a) related to the 2016 United States presidential election, which consists of abusive and discriminatory messages against a person or minority group.
Task IV: Sarcasm. Sarcastic content usually has an incongruity between the intended meaning and the way it is expressed. These are generally used to express dissatisfaction or to veil insult through humor (Chauhan et al., 2020). Given that sarcastic elements in multimodal memes often stem from incongruities between the meme’s image and text, our interest lies in investigating the capacity of LMMs for Multimodal Understanding to detect sarcasm within multimodal inconsistent contexts. To delve into this investigation, we curated meme samples from the MSD dataset (Cai et al., 2019), a compilation of authentic image-text pairs sourced from English Twitter.
Task V: Harmfulness. Harmful memes are defined as multimodal units consisting of an image and embedded text that have the potential to cause harm to an individual, an organization, a community, or society in general (Sharma et al., 2022). While our research endeavors to explore the dark side of memes, specifically targeting hateful and offensive content, it’s crucial to note that ‘harmful’ encompasses a broader spectrum than ‘offensive’ or ‘hateful’. Offensive and hateful memes fall within the realm of harmful content, yet not all harmful memes can be classified as offensive or hateful (Pramanick et al., 2021a). Thus we are also interested in exploring whether LMMs could understand the coarser-grained multimodal meme terms that transcend the scope of just being hateful or offensive. To investigate this, we employed two datasets, namely Harm-C (Pramanick et al., 2021a) and Harm-P (Pramanick et al., 2021b). These datasets encapsulate instances of harmful content within memes, often veiled and necessitating nuanced assessment to discern their potential impact.
Table 1: GOAT-Bench Sources, Scope, and Data Distribution.
Term Scope Sources #Distribution of Labels #Total
Hatefulness Target individuals based on aspects like ethnicity, gender, and disability, with dehumanizing speech and mockery of hate crimes. FHM Hateful 750 2,000
Non-hateful 1,250
Misogyny Target women with sexism or hate, involving shaming, stereotyping, objectification, or violence. MAMI Misogynistic 500 1,000
Non-misogynistic 500
Offensiveness Provoke, offend, or embarrass individuals, regardless of specific group targeting. MultiOFF Offensive 305 743
Non-offensive 438
Sarcasm Feature irony and humor to subtly express dissatisfaction or mock, often through incongruous text and image. MSD Sarcastic 910 1,820
Non-sarcastic 910
Harmfulness Cause harm to individuals, organizations, or society, beyond just being offensive or hateful. Harm-C, Harm-P Harmful 444 1,063
Non-harmful 619
2.2 Annotation and Quality Assurance
We employed binary classification for meme annotations across the five intertwined tasks: Task I: Hateful or Non-hateful; Task II: Misogynistic or Non-misogynistic; Task III: Offensive or Non-offensive; Task IV: Sarcastic or Non-sarcastic; and Task V: Harmful or Non-harmful. Please note that concerning the meme samples selected from the Harm-C and Harm-P datasets, which were initially labeled into three classes: ‘very harmful’, ‘somewhat harmful’, and ‘not harmful’, we merged the ‘very harmful’ and ‘somewhat harmful’ categories into ‘Harmful’ memes, while considering the ‘not harmful’ memes as ‘Non-harmful’ memes.
To ensure consistency with previous research, our initial annotation process involved referencing the labels from the original raw meme datasets, which were initially assigned based on hashtags in the original posts or by moderators. To reinforce the annotation quality, one professional linguistic annotator re-evaluated the data. Ultimately, we selected samples where the annotations aligned consistently between the hashtags and the human annotator.
3 Methodology
3.1 Problem Formulation
We define a meme $M=\{\mathcal{I},\mathcal{T}\}$ is a tuple representing an image $\mathcal{I}$ that is associated with a text sequence $\mathcal{T}$. In this work, to unravel LMMs’ awareness of social abuse through memes, we formulate each task as a natural language generation paradigm, where LMM take the text $\mathcal{T}$ and image $\mathcal{I}$ as the input and generates a textual output corresponding to the label $y$ to distinctly portraying whether the meme demonstrates abusive content within each specified task.
3.2 Models
To provide an exhaustive perspective on the current state of large multimodal models within the context of meme-based abuse, we conducted evaluations on 11 accessible LMMs. These models have undergone various stages of development, including pretraining, instruction tuning, and reinforcement learning from human feedback (Ouyang et al., 2022). Our selection encompasses a range of models from diverse organizations, differing in their size and complexity, which allows for a thorough understanding of the capabilities and limitations of LLMs in handling meme-based abusive content, as shown in Table 2.
For the LMMs with Pretraining, we utilized OpenFlamingo (Awadalla et al., 2023), CogVLM (Wang et al., 2023a), and Fuyu (Bavishi et al., 2023). Notably, CogVLM underwent further refinement via instruction tuning. However, it remains unclear whether Fuyu was subjected to supervised fine-tuning or reinforcement learning. In the case of LLMs solely enhanced through Instruction Tuning, our selection included InstructBLIP (Dai et al., 2023), LLaVA-1.5 (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), MiniGPT-V2 (Chen et al., 2023), MMGPT (Gong et al., 2023), mPLUG-Owl (Ye et al., 2023), and Qwen-VL (Bai et al., 2023). Additionally, GPT-4V (OpenAI, 2023), the state-of-the-art LMM was also considered in our study. Except that GPT-4V is the closed-source LMM, all the other models are open-source.
3.3 Prompting and Standards
[…]
Given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], is this meme [*]?
This deductive approach in a binary manner is designed to yield more precise and task-specific evaluations from the LMMs, facilitating a clearer understanding of their capability to identify potentially abusive usage of memes. For those still ambiguous outputs, we consider them indicative of the model’s inability to make a nuanced safety judgment regarding the meme. Notably, as different LLMs vary in their adherence to directive outputs, we adapt our prompt for each model based on the specific context to achieve controlled outputs. Moreover, we further manually draft rules with regex for filtering, to ensure aligning the model’s responses with one of the two categories in our binary classification standard, thereby facilitating a more structured and effective evaluation process.
4 Experiments
In the preceding sections, we offered an in-depth account of the structuring of the GOAT-Bench and the types of LMMs employed. This section will commence with an overview of the performance exhibited by various LMMs on the GOAT-Bench, spanning diverse settings and tasks, complemented by a thorough analysis.
4.1 Experimental Setup
We conduct extensive experiments on the GOAT-Bench to evaluate a total of 11 LMMs:
  • GPT-4V (OpenAI, 2023): The fourth iteration featuring visual capabilities within the GPT series (Radford et al., 2018), is a proprietary product of OpenAI. It has been enhanced through stages of pretraining, instruction tuning, and RLHF. We specifically utilize the “gpt-4-vision-preview” version.
  • CogVLM (Wang et al., 2023a): An open-source LMM to bridge the gap between frozen pre-trained language models and image encoders, integrating a trainable visual expert module within its attention and feed-forward layers. We specifically utilize the “cogvlm-chat” version.
  • LLaVA-1.5 (Liu et al., 2023a): An enhanced version of LLaVA that combines a vision encoder CLIP (Radford et al., 2021) with an LLM LLaMA (Touvron et al., 2023a), designed for general visual and language understanding, and developed through instruction tuning using language-image data generated by GPT-4. We specifically utilize the “llava-v1.5-13b” version.
  • InstructBLIP (Dai et al., 2023): An exploration of instruction tuning techniques in the domain of image-text multimodal large models, demonstrating the applicability of instruction tuning in the multimodal field. We specifically built the model based on the frozen “vicuna-13b-v1.1” Vicuna (Chiang et al., 2023) version with the LAVIS library (Li et al., 2022).
  • MiniGPT-4 (Zhu et al., 2023): An open-source LMM that aligns a frozen visual encoder with a frozen Vicuna, using one projection layer. We specifically utilize the “minigpt4-vicuna-13b” version.
  • Qwen-VL (Bai et al., 2023): An open-source LMM to enhance the LLM foundation with visual capabilities by adding a language-aligned visual encoder and a position-aware adapter, enabling finer perception skills like object grounding and text reading. We specifically utilize the “qwen-vl-chat” version.
  • OpenFlamingo (Awadalla et al., 2023): An open-source replication of Flamingo models (Alayrac et al., 2022) that enhances pre-trained, frozen language models by enabling them to cross-attend to the outputs of a frozen vision encoder during the next token prediction. We specifically utilize the “openflamingo-9b-vitl-mpt7b” version.
  • MMGPT (Gong et al., 2023): A vision and language model fine-tuned from OpenFlamingo, with Low-rank Adapter (LoRA) (Hu et al., 2021) incorporated in both the gated-cross-attention and self-attention components of the language model.
  • Fuyu (Bavishi et al., 2023): A multi-modal text and image transformer, utilizing a decoder-only architecture without an image encoder, where image patches are linearly projected into the transformer’s first layer.
  • mPLUG-Owl (Ye et al., 2023): A training approach that enhances LLMs with multimodal capabilities by integrating a foundational LLM with a visual knowledge module and a visual abstractor module, using a two-stage method to align image and text. We specifically utilize the “mplug-owl-llama-7b” version.
  • MiniGPT-v2 (Chen et al., 2023): A multimodal model designed as a unified interface for a variety of vision-language tasks, utilizing unique identifiers for different tasks to enhance task-specific learning efficiency and performance with three-stage training. We specifically built the model based on the “llama-2-7b-chat” LLaMA version with the checkpoint of the online developing demo.
To ensure our results are reproducible, we set the temperature as 0 without any sampling mechanism. We use the accuracy and macro-averaged F1 score as the evaluation metrics.
Table 2: Results of different LMMs on the GOAT-Bench across different tasks. The accuracy and macro-averaged F1 score (%) are reported as the metrics. The best and second test results are in bold and underlined, respectively.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
Method (#Params) Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
GPT-4V (-) 71.70 71.28 80.80 80.79 61.78 61.34 75.88 75.08 65.91 62.96 72.17 70.29
CogVLM (17B) 61.50 60.03 60.70 55.46 48.18 46.22 50.00 33.34 63.04 63.03 56.97 51.62
LLaVA-1.5 (13B) 65.20 61.40 57.10 48.68 58.68 58.39 73.52 72.89 66.70 64.35 65.77 61.14
InstructBLIP (13B) 58.25 57.42 63.10 62.70 50.61 50.24 63.41 60.96 54.11 53.93 58.88 57.05
MiniGPT-4 (13B) 58.20 39.98 52.70 52.36 53.84 45.54 55.55 54.38 56.89 44.41 55.94 47.33
Qwen-VL (10B) 64.00 56.42 71.40 71.39 59.62 59.21 60.11 54.12 61.35 53.66 63.13 58.96
OpenFlamingo (9B) 58.65 51.78 54.70 54.42 47.64 46.60 45.49 44.64 56.89 52.83 52.92 50.05
MMGPT (9B) 37.50 27.28 49.50 33.29 29.07 22.53 49.95 33.31 32.51 24.75 40.98 28.23
Fuyu (8B) 38.55 30.22 56.70 56.62 50.34 50.20 50.93 46.45 57.78 57.72 49.10 48.24
mPLUG-Owl (7B) 58.95 58.90 65.80 62.16 46.30 41.44 54.73 43.27 63.43 63.35 58.12 53.82
MiniGPT-v2 (7B) 57.35 57.27 59.00 52.00 51.82 49.05 65.55 62.63 57.38 57.07 59.24 55.60
4.2 Main Results
In Table 2, we showcase the average outcomes of 11 accessible LMMs tested in the zero-shot setting on the GOAT-Bench, each executed thrice. Based on the results in Table 2, we draw the following observations:
For the overall performance, GPT-4V, enhanced with RLHF, stands out as the preeminent model, demonstrating considerable robustness across a spectrum of tasks with the highest overall 72.17% accuracy and 70.29% macro-averaged F1 score. Instruction-tuned LMMs such as LLaVA-1.5 and Qwen-VL also deliver competitive results. These findings suggest that techniques like instruction tuning and RLHF are effective in refining the models’ alignment with human values, consequently augmenting their precision in addressing safety concerns associated with memes.
In terms of task-specific performance, LLaVA-1.5 falls behind GPT-4V but also exhibits a strong capability for detecting hatefulness, with both models obtaining notable scores when compared with other LMMs. For discerning offensiveness, GPT-4V, LLaVA-1.5, and Qwen-VL all present formidable and closely matched outcomes. In misogyny detection, GPT-4V and Qwen-VL are distinguished by their considerable accuracy and macro-averaged F1 scores, while LLaVA-1.5 appears to have a reduced performance in this aspect. Despite its lower sensitivity to sexism, LLaVA-1.5, along with GPT-4V, leads in identifying sarcasm and consistently outperforms others in recognizing harmfulness, surpassing even GPT-4V in this regard.
The results reflect a significant variation in model performance, indicating that certain models like GPT-4V, LLaVA-1.5 and Qwen-VL adapt relatively better to the complexity of multimodal meme tasks compared to others. Models with a higher number of parameters do not always guarantee better performance across all tasks, as seen with CogVLM and MMGPT. These insights highlight the intricacies involved in multimodal meme understanding, and the need for continued advancements in LMM safety and human alignment to enhance performance on complex tasks such as those presented by the GOAT-Bench.
4.3 Effect of Zero-Shot CoT Prompts
Table 3: Results of different LMMs with zero-shot CoT prompts on the GOAT-Bench across different tasks.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
Model (#Params) Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
GPT-4V (-) 71.90 71.37 81.60 81.57 61.37 60.99 77.75 77.26 66.90 64.97 72.99 71.23
CogVLM (17B) 63.30 58.70 63.20 59.77 50.87 50.17 50.44 34.30 58.18 57.74 57.54 52.14
LLaVA-1.5 (13B) 65.75 60.70 56.90 48.24 60.16 59.73 67.91 66.72 66.40 63.65 64.49 59.81
InstructBLIP (13B) 57.55 56.91 58.00 53.41 50.07 49.14 61.21 57.12 55.00 54.70 57.38 54.25
MiniGPT-4 (13B) 56.15 42.70 53.70 53.63 48.05 45.69 57.36 56.45 54.01 53.07 54.86 50.31
Qwen-VL (10B) 63.85 53.57 69.80 69.71 60.30 59.08 53.46 41.93 59.46 45.70 60.79 54.00
OpenFlamingo (9B) 51.80 51.69 52.00 49.61 44.15 38.89 50.33 34.45 53.82 53.82 50.89 45.69
MMGPT (9B) 37.45 27.31 50.10 33.56 36.74 26.87 49.95 33.59 40.44 28.79 43.19 30.02
Fuyu (8B) 41.75 37.22 50.90 46.62 46.03 42.25 41.59 41.24 50.35 49.69 44.95 43.40
LLaVA-1.5 (7B) 53.25 50.79 49.40 32.76 41.59 30.48 50.11 33.58 44.00 34.52 49.02 36.43
mPLUG-Owl (7B) 51.00 50.36 61.50 55.88 44.68 37.76 53.90 41.69 57.28 56.75 53.68 48.49
MiniGPT-v2 (7B) 60.20 59.65 58.50 51.28 52.09 50.08 66.16 63.56 58.28 58.03 60.36 56.52
[…]
Given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], let’s think step by step, is this meme [*]?
As indicated in Table 3, GPT-4V, when prompted with CoT, significantly outperforms other open-source LMMs. In comparison to Table 2, we can observe that for each individual task, approximately 5 to 6 LMMs show improvement with CoT prompts, while an equal number experience performance degradation. A more visual comparison in Figure 3 reveals that models such as GPT-4V, CogVLM, MiniGPT-4, MMGPT, and MiniGPT-v2 benefit overall from CoT prompts. In contrast, LLaVA-1.5, InstructBLIP, Qwen-VL, OpenFlamingo, Fuyu, and mPLUG-Owl demonstrate reduced performance compared to the setting without CoT. Thus, the effectiveness of CoT in enhancing performance is contingent on the specific model and task type under consideration for our GOAT-Bench.
Figure 3: The comparison among the overall macro-averaged F1 scores (%) of different LMMs with CoT prompts on the GOAT-Bench across different tasks.
4.4 Effect of Few-Shot In-context Learning
[…]
You are a detector trained to identify meme-based social abuse. Here are some examples: **example1** Given a meme: [example1], is this meme [*]? Answer: Yes. **example2** Given a meme: [example2], is this meme [*]? Answer: No. Based on the meme examples, now given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], is this meme [*]?
It is noteworthy that 1) the LMM’s performance significantly declines with an increase in the number of few-shot, in-context learning examples. This decline may be attributed to the LMMs’ difficulty in abstracting generalized reasoning patterns from the complex, multimodal thinking processes inherent in a limited set of meme examples. This suggests that while the LLaVA can handle straightforward tasks with minimal data, their ability to scale up reasoning capabilities in more intricate scenarios is limited. 2) Simultaneously, except for the other four meme tasks, LLaVA only manages to glean some performance improvements in the finer-grained Misogyny task from a finite set of ICL demonstrations, yet this improvement is marked by a degree of fluctuation due to the unpredictable nature of meme content and context. This indicates that while LLaVA can adapt to specific, narrowly defined tasks, its performance is not consistently reliable across broader or more variable contexts. 3) Overall, the in-context learning prompting strategy does not yield a consistent upward trend in performance or enhance the model’s robustness. This reflects the formidable challenge meme-based social abuse poses to the LMM paradigm, underscoring the intricate complexities and nuanced understandings required to effectively navigate and interpret such content in future multimodal research about large language models. We speculate that the primary reason for these challenges is likely due to the model’s lack of sufficient training on multi-image scenarios. This insufficiency means that LMMs are not fully equipped to handle the diverse and layered contexts presented by memes, which often combine text, images, and cultural references in ways that are difficult to parse without extensive, specialized training. Addressing this gap will be crucial for future developments in multimodal learning, as it highlights the need for training data that better reflects the complexity of real-world applications.
Table 4: Results of LLaVA with ICL prompts on the GOAT-Bench across different tasks.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
#Examples Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
0-shot 65.20 61.40 57.10 48.68 58.68 58.39 73.52 72.89 66.70 64.35 65.77 61.14
1-shot 62.70 60.02 68.70 68.05 46.30 40.50 63.08 62.98 47.47 44.52 59.43 55.21
2-shot 63.00 42.19 58.40 52.74 59.49 52.44 62.47 60.72 56.29 56.27 60.69 52.87
3-shot 62.80 40.06 55.20 45.71 60.97 48.37 58.57 52.39 59.56 55.38 59.77 48.38
4-shot 63.20 50.06 67.60 66.62 49.66 48.05 52.69 44.68 53.02 52.49 57.83 52.38
5-shot 63.00 55.96 69.10 69.08 46.97 42.52 51.92 44.57 53.22 52.40 57.51 52.91
6-shot 40.60 35.11 52.10 51.31 39.57 35.89 48.30 39.23 41.33 40.71 44.45 40.45
4.5 Case Study
To better understand the behavior of the best-performed model GPT-4V and facilitate future studies on GOAT-Bench, we provide case studies on both the wrongly and the correctly predicted memes by GPT-4V. To make the analysis more intuitive, we conduct abductive reasoning for each example to obtain the output explanation by GPT-4V for the prediction, as shown in Figure 4-6.
Figure 4: Hateful example of wrongly predicted memes by GPT-4V with the explanation.
Figure 5: Non-hateful example of wrongly predicted memes by GPT-4V with the explanation.
Figure 6: Misogynistic example of correctly predicted memes by GPT-4V with the explanation.
From the perspectives of both the prediction result and the generated explanation, we can analyze that: 1) It is evident from Figure 4 that the best-performed model GPT-4V failed to detect the implicit connotation within the association of the term ‘gas’ in the meme text with the historical figure depicted as Adolf Hitler. This oversight for the wrongly predicted example is significant as it underplays the severity of historical atrocities. Such a juxtaposition in a seemingly trivial context might be construed as insensitive and hate, potentially minimizing the enormity of the suffering experienced by the victims of that era. 2) In terms of Figure 5, we exemplify the false positive that one non-hateful example is incorrectly perceived as a hateful meme. The model’s output explanation suggests that it mistakenly associates the image and text with xenophobic or inflammatory sentiment, due to its exposure to similar content targeting race or religion, especially involving Muslim communities. This misclassification highlights a potential bias in the LMM, where content involving certain groups is more likely to be flagged as hateful. Moreover, the explanation provided by the model lacks coherence and precision, as it overgeneralizes and introduces unnecessary context. The rationale here is not well-aligned with the actual content of the meme, indicating that the model may rely too heavily on superficial features such as the appearance of the individual or the use of certain keywords. 3) As demonstrated in Figure 6, we observe that despite the model arriving at the correct classification, the explanation provided fails to capture the underlying reasoning effectively. It is imperative for a human checker to discern that the representation of ‘respect’ as a firearm implies a menacing undertone towards the women depicted as prostrate. This interpretation, which is vital for a comprehensive understanding of the meme’s implications, is not adequately articulated in the explanation generated by GPT-4V.
If the LMM exhibits bias towards content from certain groups, it could result in unfair treatment of those groups in real-world applications. For example, if the LMM misclassifies content from specific groups as hateful or offensive during online content moderation, it could lead to unjust removal or censorship of their content, exacerbating information inequality. Additionally, the model’s explanations might reinforce bias in some cases by incorporating unnecessary social context into its reasoning, further embedding potential prejudice into the interpretation. Our proposed GOAT-Bench underscores the pressing need for safety insights into the emerging LMMs to be closely aligned with human values in future research. This alignment is critical to ensure that LMM-generated responses avoid causing harm or discomfort.
4.6 Effect of Supervised Fine-Tuning
To investigate how to synchronize the generative and detection capabilities of LMMs with their emergent abilities regarding meme-based social abuse, and how to minimize the need for extensive human intervention, we explore an innovative instruction-tuning strategy, SelfAlign, for improving LMMs’ safety insights on meme-based social abuse. This approach enables LMMs to self-align iteratively with minimal active human engagement for improving the safety awareness of meme-based social abuse. Specifically, we first conduct abductive reasoning to prompt the LMMs, to generate a rationale that elicits the reasoning knowledge about how to infer the golden-truth harmfulness label of a training data sample. Then we conduct instruction tuning, by concatenating the golden-truth harmfulness label and the generated rationale as the targeted output of LMMs for training purposes. To this end, we have meticulously curated a dataset comprising 2,000 samples specifically for instruction tuning across five meme-related tasks, allocating 500 samples to each task for precise tuning efforts. We select the representative open-source LMM, LLaVA, as our foundation model for parameter-efficient tuning (Hu et al., 2021), specifically the “llava-v1.5-7b” version to be instruction-tuned on two NVIDIA A100 80GiB GPUs.
For abductive reasoning, we curate a template $p$ that consists of a triplet $\{y,{\mathcal{I}},\mathcal{T}\}$ as observed attributes, to prompt the LMMs to generate a rationale $r$ that elicits the reasoning knowledge about how to infer the label $y$ based on the interplay of the meme text $\mathcal{T}$ and the image ${\mathcal{I}}$. Specifically, we design the prompt $p$ as:
Given the text [$\mathcal{T}$], which is accompanied by the image; and a harmfulness label [$y$], please give me a streamlined rationale associated with the meme, without explicitly indicating the label, for how it is reasoned as [$y$].
By explicitly incorporating the ground-truth label within the prompt’s observed attributes, we can effectively mitigate the inherent biases and variations present in LMMs, which facilitates the activation of rich contextual background knowledge through abductive reasoning, grounded in the truth, thereby naturally and implicitly filtering out invalid rationales.
For instruction tuning, we propose two variants for detecting meme-based social abuse: 1) LLaVA-Label: we simply utilize the golden label $y$ as the targeted output for supervised fine-tuning, with the input prompt as designed in §3.3; 2) LLaVA-SelfAlign: we concatenate the classification ground truth $y$ and the generated rationale $r$ as the targeted output for supervised fine-tuning, with the input prompt as designed in §3.3.
Table 5: Results of LLaVA with zero-shot prompt and supervised fine-tuning on the GOAT-Bench across different tasks.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
Model (#Params) Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
LLaVA-1.5 (13B) 65.20 61.40 57.10 48.68 58.68 58.39 73.52 72.89 66.70 64.35 65.77 61.14
LLaVA-1.5 (7B) 52.80 52.58 50.00 33.51 55.85 55.84 50.00 33.34 59.46 57.95 53.02 46.64
LLaVA-CoT (7B) 53.25 50.79 49.40 32.76 41.59 30.48 50.11 33.58 44.00 34.52 49.02 36.43
LLaVA-Label (7B) 63.50 53.70 67.30 65.10 46.03 40.09 73.02 72.97 71.56 71.54 66.02 60.68
LLaVA-SelfAlign (7B) 63.05 54.06 75.10 74.66 53.30 52.33 75.16 74.80 66.11 65.41 67.59 64.25
The results from Table 5 indicate that the LLaVA models with different parameter sizes and prompting/training strategies exhibit varying performance across multiple tasks, including Hatefulness, Misogyny, Offensiveness, Sarcasm, and Harmfulness. 1) Parameter Size Impact: The LLaVA-1.5 (13B) generally outperforms the LLaVA-1.5 (7B) across all tasks, suggesting that a larger parameter count contributes positively to model performance. This is evident in the overall accuracy and F1 scores, where the 13B version leads by a notable margin. 2) Fine-Tuning Advantage: The LLaVA-Label (7B) and LLaVA-SelfAlign (7B) surpass both the zero-shot version LLaVA-1.5 (7B) and achieve competitive performance compared with LLaVA-1.5 (13B) in almost all tasks, especially in Misogyny and Sarcasm, where the improvement is significant. This underscores the effectiveness of our proposed supervised fine-tuning variants tailored for meme-based social abuse, which seems to enhance model performance even with a smaller parameter size. 3) Task-Specific Observations: In Hatefulness, Misogyny, and Harmfulness, the accuracy and F1 scores improve with supervised fine-tuning, highlighting the advantage of aligning the model’s predictions with the labeled data. For Offensiveness, the LLaVA-1.5 (7B) has higher accuracy and F1 score than its 13B counterpart. Sarcasm detection benefits the most from fine-tuning, with LLaVA-SelfAlign (7B) achieving the highest accuracy and F1 score. This may indicate that sarcasm detection is more sensitive to model training approaches and can greatly benefit from human-instruction-free fine-tuning. 4) Overall Performance: Overall, the self-align instruction-tuning method with the 7B model achieves the best balance between recall and precision, as evidenced by its superior F1 score compared to its LLaVA-Label counterpart and both the 13B and 7B zero-shot models. These findings suggest that while larger models have inherent advantages, the fine-tuning process, particularly with a targeted and refined training approach like our proposed SelfAlign, can significantly enhance a model’s ability to understand and predict complex multimodal tasks related to meme-based social abuse.
4.7 Sense of Humor
Figure 7: The performance comparison among the accuracy and macro-averaged F1 scores (%) of different LMMs on humor detection.
Considering that social abuse in memes often masquerades under the guise of humor, we further dive into whether LMMs can discern humor akin to human perception. To this effect, we collected 1,152 code-mixed memes consisting of the same humorous and non-humorous samples originally labeled by Mishra et al. (2023). These meme samples serve to evaluate the proficiency of LMMs in detecting humor in the context of multimodal and multilingual memes. Figure 7 demonstrates a comparative analysis of various LMMs in humor detection, focusing on accuracy and macro-averaged F1 score. It is observed that models employing CoT prompting typically demonstrate enhanced performance, with OpenFlamingo showing particularly notable improvements. The Fuyu and LLaVA-1.5 models excel in both accuracy and F1 score, irrespective of CoT usage. However, the impact of CoT prompting is not uniformly beneficial across all models. For instance, while OpenFlamingo benefits significantly from CoT, MiniGPT-4 experiences a performance decline, suggesting that the effectiveness of CoT prompting varies depending on the specific model architecture and task nature. Overall, even the best-performed models hover around a macro-averaged F1 score of 50%, indicating that LMMs are not highly effective humor detectors. Given the close relationship between humor and memes, enhancing a model’s understanding of humor could significantly improve its ability to perceive meme-based social abuse, making these two aspects mutually reinforcing.
4.8 Trend with Emerging LMMs
Table 6: Results of three representative emerging LMMs after July 2024 on the GOAT-Bench across different tasks. The accuracy and macro-averaged F1 score (%) are reported as the metrics.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
Model (#Params) Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
GPT-4o (-) 71.70 68.95 81.20 81.12 62.13 61.16 78.79 78.40 66.01 64.07 73.10 70.74
w/ zero-shot CoT 72.00 68.58 82.66 82.52 62.26 61.22 79.12 78.70 67.00 64.20 73.67 71.04
[…]
w/ zero-shot CoT 72.28 70.10 79.19 79.06 53.10 52.80 68.89 66.46 54.96 54.61 67.46 64.61
[…]
w/ zero-shot CoT 70.30 69.00 80.00 79.54 59.08 59.08 60.33 53.98 66.90 66.06 67.22 65.53
[…]
w/ zero-shot CoT 70.30 63.13 69.70 68.17 59.49 55.36 58.52 52.81 62.14 59.24 64.45 59.74
[…]
w/ zero-shot CoT 68.90 68.02 73.40 72.49 63.26 57.67 67.09 66.47 59.27 50.23 66.90 62.98
Since most of the LMMs evaluated in Table 2 were released prior to March 2024 when this work was originally submitted for review, we further incorporated five representative emerging LMMs (i.e., GPT-4o, Claude-3-5-Sonnet, GPT-4o-mini, MiniCPM-V (Yao et al., 2024), and Qwen2-VL (Wang et al., 2024)) released after July 2024 to strengthen the robustness of our findings:
  • GPT-4o: the latest flagship model developed by OpenAI, designed for real-time reasoning across audio, visual, and textual inputs. It excels in understanding both vision and audio, offering significant improvements over previous models in these areas. We specifically utilize the “gpt-4o-2024-11-20” version.
  • Claude-3-5-Sonnet: The LMM developed by Anthropic with significant improvements most evident in visual reasoning tasks like interpreting charts and graphs. We specifically utilize the “claude-3-5-sonnet-20241022” version.
  • GPT-4o-mini: The mini version of GPT-4o. We utilize the “gpt-4o-mini-2024-07-18” version.
  • MiniCPM-V: The latest and most capable model in the MiniCPM-V series developed by OpenBMB, achieves an average score of 65.2 on the latest version of OpenCompass, a comprehensive evaluation over 8 popular benchmarks. We utilize the “MiniCPM-V-2_6” version.
  • Qwen2-VL: The latest addition to the vision-language models in the Qwen series, building upon the capabilities of Qwen-VL. We specifically utilize the “Qwen2-VL-7B-Instruct” version.
For the two open-source LMMs, note that Qwen2-VL and MiniCPM-V differ significantly in their approaches to visual encoding and position embeddings. Qwen2-VL employs a "Naive Dynamic Resolution" mechanism to process images of varying resolutions, mapping them into dynamic visual tokens, and enhances multimodal understanding with Multimodal Rotational Position Embeddings (M-ROPE) to capture 1D, 2D, and 3D positional information. In contrast, MiniCPM-V uses a Perceiver Resampler to compress image representations into a fixed number of tokens for efficiency, focusing on lightweight deployment without advanced multimodal position embedding mechanisms.
As shown in Table 6, we can find that GPT-4o performs best overall with zero-shot CoT, achieving the highest dominant metric F1 score (71.04%) among the five recent popular LMMs. Except for GPT-4o, Claude-3-5-Sonnet leads in understanding the hatefulness with CoT, and GPT-4o-mini leads in understanding the harmfulness while sarcasm remains the most challenging task for GPT-4o-mini. Compared with Claude-3-5-Sonnet, MiniCPM-V and Qwen2-VL, zero-shot CoT only improves the performance of GPT-4o and GPT-4o-mini in most tasks, especially for misogynistic, offensive and sarcastic content. Each model has specific strengths, with trade-offs depending on the task, making GPT-4o the most consistent performer overall. To offer a more meaningful view of in-context learning as shown in Table 4, we further incorporate the representative LMM, Qwen2-VL, trained with in-context learning capabilities, to conduct a comprehensive assessment of in-context learning. As shown in Table 7, we can observe that it has similar trends to the results illustrated in Table 4, which highlights ICL performance of meme-based social abuse is still a challenge for the reasoning ability of open-source LMMs as the unsafe memes usually consist of safe texts and safe images.
Table 7: Results of Qwen2-VL with ICL prompts on the GOAT-Bench across different tasks.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
#Examples Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
0-shot 70.20 68.58 75.30 75.16 62.99 54.94 70.27 70.24 59.17 48.28 68.41 63.44
1-shot 69.15 63.89 70.80 69.80 62.72 53.97 50.11 33.58 62.93 56.54 62.45 55.56
2-shot 69.00 63.41 70.40 69.41 62.72 52.70 50.00 33.34 59.56 45.02 61.77 52.78
3-shot 68.30 62.82 69.00 67.46 64.06 56.25 50.11 33.58 61.74 50.01 61.88 54.02
4-shot 68.65 62.41 71.10 70.53 63.12 55.65 50.27 34.04 59.27 44.23 61.85 53.37
5-shot 67.90 61.22 70.20 69.38 63.66 56.86 50.71 35.46 59.96 46.00 61.78 53.78
6-shot 67.80 59.11 70.50 70.03 64.33 59.27 50.77 35.57 60.46 47.31 61.96 54.26
4.9 Cross-Lingual Detection Performance
Figure 8: The performance comparison among the accuracy and macro-averaged F1 scores (%) of LLaVA, GPT-4V and Qwen2-VL on cross-lingual detection of meme-based social abuse.
To provide a more nuanced understanding of LMMs’ capabilities across various languages and cultural contexts, we investigate the cross-lingual detection performance of three top-performing LMMs: LLaVA, GPT-4V (both released before Mar. 2024), and Qwen2-VL (after Jul. 2024), where LLaVA and Qwen2-VL representing the open-source LMMs, and GPT-4V representing the closed-source LMMs, respectively. We evaluate these models, pre-trained in English, on three datasets: MET-Meme (Xu et al., 2022) (Chinese), TamilMemes (Suryawanshi et al., 2020b) (Tamil), and MIMIC (Singh et al., 2024) (code-mixed Hindi-English).
Figure 8 compares the performance of three LMMs, LLaVA, GPT-4V, and Qwen2-VL, on three datasets (MET-Meme, TamilMemes, and MIMIC) using accuracy and macro-average F1 score metrics. GPT-4V consistently outperforms LLaVA and Qwen2-VL on the MET-Meme and MIMIC datasets, demonstrating superior cross-lingual detection capabilities in those contexts. For MET-Meme (Chinese), GPT-4V achieves 77.75% accuracy and 68.09% F1, while LLaVA reaches about 66.37% and 61.27%, Qwen2-VL gets about 70.72% and 65.21%, respectively. In the MIMIC dataset (code-mixed Hindi-English), GPT-4V outperforms LLaVA with approximately 50.95% accuracy and 33.76% F1, while LLaVA scores around 25.82% and 20.52%, Qwen2-VL reaches around 49.37% and 33.05%. The performance gap is most pronounced in the MIMIC dataset between GPT-4V and LLaVA, indicating GPT-4V’s stronger handling of code-mixed language scenarios. However, on the TamilMemes dataset (Tamil), LLaVA performs better with about 51.63% accuracy and 47.42% F1, compared to GPT-4V’s 44.83% and 44.32%, and Qwen2-VL’s 38.42% and 34.48%. This indicates that while GPT-4V and Qwen2-VL generally have stronger cross-lingual detection capabilities than LLaVA, LLaVA is more effective in detecting Tamil memes. Overall, all the LMMs struggle in cross-lingual scenarios, indicating that the complexity of understanding and detecting nuances in multilingual memes remains a significant challenge, highlighting the need for more advanced approaches and improvements for LMMs in handling such diverse cultural contexts.
4.10 Comparison with Non-LMM Baseline
Table 8: Results of the CLIP-based model on the GOAT-Bench across different tasks.
Task Hatefulness Misogyny Offensiveness Sarcasm Harmfulness Overall
#Examples Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
0-shot 50.25 48.75 58.30 57.42 45.36 44.94 60.60 57.15 51.14 49.50 53.90 51.55
1-shot 43.95 43.40 51.80 42.95 53.98 53.09 51.54 37.02 58.37 57.77 50.66 46.85
2-shot 50.80 50.71 50.10 41.10 54.80 54.47 59.95 53.95 60.26 60.21 55.17 52.09
3-shot 55.10 54.44 50.70 36.64 56.99 56.80 66.54 63.94 61.74 61.73 58.86 54.71
4-shot 51.00 51.00 54.90 45.40 57.96 57.07 64.56 62.02 64.22 64.15 58.21 55.93
5-shot 52.95 52.78 54.00 43.68 58.53 57.48 64.07 60.74 65.91 65.87 58.87 56.11
6-shot 55.25 54.51 52.50 40.24 57.18 57.10 65.05 62.82 66.50 66.30 59.55 56.19
We conduct more experiments by applying the CLIP-based model (Radford et al., 2021) as the non-LMM baseline on the GOAT-Bench in zero-shot and few-shot settings. As depicted in Table 8, the CLIP-based model’s performance varies across tasks, with the best overall F1 score (56.19%) achieved in the 6-shot setting. Sarcasm detection peaks in the 3-shot setting with an F1 of 63.94%, while offensiveness sees its highest F1 score (57.48%) at 5-shot. For hatefulness, the model achieves its best F1 score (54.51%) in the 6-shot setting. Misogyny detection performs best in 0-shot (57.42%), highlighting the model’s initial capability without additional examples. Harmfulness detection improves steadily, with the highest F1 score (66.30%) in 6-shot. Different from the findings in Table 4, increasing shots generally improves the performance of the non-LMM baselines, though the effect varies by task. This observation reaffirms that the ICL ability of existing LMMs for meme-based social abuse needs to be improved, and the emerging LMMs are more promising to exploit in the future.
4.11 Discussion
For the open-source LMMs, test set leakage is not a concern, as the literature explicitly delineates the datasets and instruction-tuning procedures employed in their training, none of which encompass the multimodal meme data utilized in our GOAT-Bench. However, we cannot fully guarantee the exclusion of potential data leakage with GPT-4V, as its internal workings remain opaque. Nevertheless, as evidenced by the results in Table 2-3, where all LMMs were evaluated directly on the GOAT-Bench, the absence of significant test set leakage is implied. This is inferred from the fact that direct application of the LMMs did not yield disproportionately high performance, which would be expected if the models were benefiting from test set leakage.
5 Related Work
5.1 LLMs and LMMs
Recently, LLMs have showcased exceptional prowess across a diverse array of tasks. Leading technology companies have made significant advancements in crafting highly proficient LLMs, like OpenAI’s GPT-3 (Brown et al., 2020) and GPT-4 (OpenAI, 2023), Google’s PaLM (Chowdhery et al., 2022) and Gemini (Team et al., 2023), DeepMind’s Gopher (Rae et al., 2022), as well as Anthropic’s Claude. However, these models are proprietary, accessible only through specific APIs or may not be accessible at all. In contrast, the AI community has witnessed the emergence of various open-source LLMs, sharing model weights with the public. Notable contributions include EleutherAI’s GPT-NeoX-20B (Black et al., 2022), Google’s UL2-20B (Tay et al., 2023), Tsinghua University’s GLM-130B (Zeng et al., 2023), and Meta’s OPT (Zhang et al., 2022) and LLaMA1&2 (Touvron et al., 2023a, b). Additionally, the performance of these open-source LLMs are further enhanced by a large amount of alignment works (Wang et al., 2023b; Xu et al., 2023; Luo et al., 2023b, a; Mukherjee et al., 2023; Zhou et al., 2023; Li et al., 2023b).
LLMs represent a significant leap forward, enabling comprehensive comprehension not only of textual data but also of visual information within a unified framework. Various approaches have emerged in constructing these advanced LMMs, exhibiting adeptness in comprehending both images and text sans the necessity for specific task-oriented training. For instance, Flamingo (Alayrac et al., 2022) boasts multifaceted capabilities by amalgamating a fixed vision module with a robust language model tailored for unified image and text understanding. Conversely, PaLM-E (Driess et al., 2023) seamlessly integrates visual cues into the highly parameterized PaLM model, comprising 520 billion parameters, thereby showcasing efficacy in real-world applications.
Recent endeavors by researchers (Yang et al., 2023) have culminated in the development of high-caliber, diverse multimodal datasets stemming from GPT-4 and GPT-4V (OpenAI, 2023), tailored for fine-tuning open-source large multimodal models. Notable examples encompass LLaVA (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), mPLUG-Owl (Ye et al., 2023), InstructBLIP (Dai et al., 2023), and other pioneering contributions (Bai et al., 2023; Wang et al., 2023a; Gong et al., 2023; Team et al., 2023; Bavishi et al., 2023).
5.2 Benchmarks for LMMs
Traditional multimodal benchmarks have historically focused on specific skills such as visual recognition (Goyal et al., 2017), image description (Agrawal et al., 2019), and visual commonsense reasoning (Zellers et al., 2019), among others. However, the emergence of advanced LMMs has generated a growing need for new benchmarks. The robust zero-shot capabilities of LMMs often surpass the capabilities evaluated by traditional assessment metrics. This limitation is compounded by their inadequacy in accurately matching given answers, leading to significant issues in robustness. Addressing these challenges, research communities have introduced a diverse array of benchmarks, including MME (Fu et al., 2023), MMBench (Liu et al., 2023b), MM-Vet (Yu et al., 2023), SEED-Bench (Li et al., 2023a), and LAMMBench (Yin et al., 2023), etc. These benchmarks facilitate structured evaluations of intricate multimodal tasks, aiming to mitigate the limitations inherent in traditional assessment methodologies. Different from these benchmarks, GOAT-Bench is specifically developed to systematically and comprehensively gauge model safety and alignment with human values, particularly in the context of social responsibility as it pertains to meme-based social abuse. While the FHM dataset (Kiela et al., 2020) has been initially used to evaluate some prior LMMs (Chen et al., 2023; Dai et al., 2023), there is a need for a more comprehensive testbed beyond the scope of hatefulness. Such a testbed would enable a deeper investigation into the safety awareness of LMMs across a broader spectrum of task types related to online abuse of memes.
6 Ethics and Broader Impact
The aim of this research focuses on the safety issue related to LMMs, to curb the dissemination of abusive memes and protect individuals from exposure to bias, racial, and gender-based discrimination. However, we acknowledge the risk that malicious actors might attempt to reverse-engineer memes that could evade detection by AI systems trained on LMMs. We vehemently discourage and denounce such practices, and emphasize that human moderation is essential to prevent such occurrences. Aware of the potential psychological impact on those evaluating abusive content, we have instituted protective measures for our human evaluators, including: 1) explicit consent regarding exposure to potentially abusive content, 2) a cap on weekly evaluations to manage exposure and advocate for reasonable daily workloads, and 3) recommendations to discontinue their review should they experience distress. We also conduct regular well-being checks to monitor their mental health. Additionally, the use of Facebook’s meme dataset necessitates adherence to Facebook’s terms of use; our use of these memes complies with these terms. Respecting Facebook’s licenses on the memes, the GOAT-Bench only contains the annotated text for the Facebook data, but not the hateful memes; users will have to download the memes from the Facebook Hateful Meme challenge separately. It is important to note that all data organized are restricted to meme content and do not include any personal user data.
7 Conclusion
In this study, we aim to investigate the safety insight of LMMs by examining the ability of LMMs to recognize social meme abuse across a spectrum of task categories. For this purpose, we have developed the GOAT-Bench, a comprehensive testbed consisting of 6,626 memes, spanning five tasks of varied complexity. Our evaluation of various LMMs using different prompting and/or instruction-tuning methods, including those with CoT, ICL prompts, or SelfAlign fine-tuning, on the GOAT-Bench reveals that these models still exhibit limitations in accurately addressing meme-based social abuse tasks. Enhancing the safety awareness and reasoning capabilities of LMMs in dealing with the intricate and nuanced tasks presented by multimodal memes is an ongoing challenge. We hope our research and the GOAT-Bench will serve as a foundation for future endeavors in this field, spurring the development of more advanced and capable large multimodal models.